OpenStudy (anonymous):

A contractor wants to use 34 feet of molding cut in 3 pieces to trim sides and top of a garage door. The long piece is 1.5 feet longer than three times the length of each shorter piece. Find the length of each piece.

OpenStudy (amistre64):

2s+l = 34 l = 34-2s l = 3s 3s = 34-2s 5s = 34 s = 34/5

OpenStudy (amistre64):

that was close but i forgot a part lol

OpenStudy (amistre64):

l = 3s + 1.5 ...

OpenStudy (amistre64):

34 - 2s -1.5 -3s --------- 32.5 -5s = 0 s = 32.5/5

OpenStudy (amistre64):

shorts are 6.5 long is 21